OpenStudy (anonymous):

Which equation can be used to find of 27% of 225? 27/100 = x/225 27/100 = x/100 100/225 = x/27 x/27 = 100/225

OpenStudy (aroub):

27/100 = x/225

jhonyy9 (jhonyy9):

225 --- 100% x ---- 27% x=27*225/100